OpenStudy (lgbasallote):

Twelve people are given two identical speakers to listen for differences, if any. Suppose that these people answered by guessing only. Find the probability that three people claim to have heard a difference between the two speakers.

OpenStudy (lgbasallote):

i suppose binomial distribution is applicable here...since there are only two possible results (difference or no difference)

OpenStudy (lgbasallote):

so n = 12 x = 3 \[\huge \implies _{12}C_3 (\frac 14)^3 (\frac 34)^9\] yes?

OpenStudy (lgbasallote):

which gives 0.2581

OpenStudy (lgbasallote):

so the probability is 25.81%

OpenStudy (lgbasallote):

thanks

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

i understood 12C3 but why 1/4 & 3/4?

OpenStudy (lgbasallote):

because of the formula \[\huge _nC_x (p)^x (q)^{n-x}\] where: n = total number of subjects x = number of people taken from that n p = probability of success q = probability of failure

OpenStudy (lgbasallote):

anyway...you get a medal for replying

OpenStudy (anonymous):

that's what i didn't get. isn't p 1/2 because you want the person to detect the difference & there are only 2 possibilities?....i'm sorry if i'm bothering you...

OpenStudy (lgbasallote):

3/12 = 1/4

OpenStudy (lgbasallote):

am i wrong?

OpenStudy (anonymous):

*smacks himself in the head* oooh i get it.

OpenStudy (lgbasallote):

hmm you may be on to something... i might be wrong

OpenStudy (lgbasallote):

what I did was wrong. thanks for making me realize (now you really deserve that medal)

OpenStudy (anonymous):

i have to agree with whatever you say because i've never seen that formula & I have no idea what "probability of success" really means....

OpenStudy (lgbasallote):

i do though....and it seems i really am wrong